# Blockchain-empowered Federated Learning: Benefits, Challenges, and Solutions

### v1

@@ 1 Introduction
Federated learning (FL) represents an emerging paradigm in the realm of privacy-preserving distributed (Machine Learning) ML. It offers a crucial technology that ensures users’ data privacy and confidentiality. Despite the convenience facilitated by the Internet, users’ information remains under the control of Internet service providers [1, 2, 3]. Artificial intelligence (AI) technology plays a prvotal role in driving the fourth industrial revolution [4]. User data serves as the foundation for training various AI models, rendering it immensely valuable [5, 6, 7]. However, service providers can easily acquire and store user data, thereby encroaching upon user privacy. Users typically have limited control over such issues [8]. In recent years, various Internet-developed countries or regions such as the European Union [9, 10], the United States [11], and Singapore [8] have introduced relevant laws or regulations to govern the use of personal privacy data and promote personal privacy protection. In addition, the emergence of FL has helped mitigate the prevalent issue of data silos, which refers to the isolation or scattering of data across different systems, departments, and organizations, making access and analysis challenging [12]. Even though the exponential growth in data volumes in the age of the internet, the data island phenomenon persists, and in some cases, has worsened. This persistence is primarily attributed to the reluctance of various organizations to share their data, driven by concerns related to privacy or competition. For example, hospitals typically refrain from sharing patient data to protect sensitive information [13].
Despite the numerous benefits of FL, several significant challenges persist that require attention. First and foremost, there is a critical need for an open and transparent incentive mechanism to encourage participation from data owners. Since the model’s training relies on the local resorces of these data owners, those issuing FL tasks must offer appropriate compensation as a reward to attract their participation in the training process. Secondly, security in the FL system remains inadequate and susceptible to vulnerabilities. Malicious nodes may deliberately provide false information, disrupting model convergence and overall training efficacy. Additionally, curious servers might infringe upon data owners’ privacy, posing additional security concerns. Lastly, the presence of a single parameter server in the FL system poses a potential bottleneck, compromising system performance and security. As more clients engage in model training, the network bandwidth and response speed of the parameter server may restrict the maximum achievable model training performance. Moreover, this central point of failure becomes a prime target for malicious attackers, making it imperative for researchers to explore novel methods for mitigating these risks and enhancing system scalability.
The integration of blockchain technology presents a promising solution to address the previously discussed shortcomings of FL. The impressive utilization of blockchain technology in Bitcoin showcases its potential to improve the security and reliability of FL systems [14]. Essentially, blockchain is a decentralized distributed ledger system [15], aligning seamlessly with the principles of FL. Blockchain technology offers several key advantages, including data transparency on the ledger [16], a consensus algorithm that ensures accurate record-keeping, and the ability to establish trust relationships in untrusted environments. At the same time, once data ois confirmed and added to the chain, it is generally immune be tampered [17]. In the FL system, where mutual trust between the parameter server and clients may not guaranteed, blockchain can play a pivotal role in fostering a trustworthy relationship between them. The emergence of the blockchain 2.0 era has enabled users to create smart contracts that operate on the blockchain [18, 19]. These smart contracts function automatically and transparently, allowing developers to incorporate various algorithms, such as security algorithms, thereby endowing the scalability of the blockchain system. Consequently, numerous studies have been conducted to explore the reasonable application of blockchain technology in FL, with the aim of overcoming its limitations and improving overall performance.
Figure 1: Main scope of this survey. Firstly, we delve into the rationale behind the integration of blockchain with the Federated Learning (FL) system and explore the methods for seamless integration. Then, we closely examine the myriad advantages that blockchain can bring to the FL system. Moving forward, our survey delves into the supplementary challenges that emerge when incorporating blockchains into the FL system, along with an exploration of the existing approaches for mitigating these challenges. Finally, we put forth recommendations for future research directions in the realm of Blockchain-empowered Federated Learning (BC-FL) systems.
This paper offers a comprehensive analysis of recent BC-FL research, commencing with an exploration of the intrinsic characteristics of blockchain and their potential advantages for FL systems. Subsequently, we delve into the challenges posed by the introduction of blockchain into FL systems and examine how current research endeavors address these challenges. Finally, we share our insights regarding prospective research avenues in the BC-FL domain. The overall structure and main scope of this survey is illustrated in Fig. 1.
The paper is structured as follows: In Section 2, we provide an introduction to the background knowledge of FL and blockchain. Section 3 elucidates the general architecture of the BC-FL system and outlines its specific benefits. Section 4 scrutinizes the additional challenges arising from the integration of blockchain into FL systems and spotlights relevant solutions. Section 5 delves into future research directions for BC-FL systems.
@@ 2.1 Federated Learning
Federated Learning (FL) was initially introduced by Google in 2016 as a privacy-preserving distributed machine learning paradigm [20]. It fundamentally represents a distributed ML approach with a specific emphasis on preserving privacy, setting it apart from conventional distributed learning methods. In the realm of FL, each client maintains ownership of their data and is dedicated to safeguarding it from potential exposure [21]. To achieve this objective, FL facilitates collaborative model training by permitting clients to locally train a model and subsequently upload it to a parameter server for aggregation [22].
2.1.1 Federated Learning Workflow
The FL process may vary depending on the specific task, but it typically adheres to a standard training procedure. An illustration of the workflow of a FL system is presented in Figure 2.
Figure 2: Workflow of a FL system. Firstly, the task publisher initializes one or more ML models along with its associated training parameters. In each iteration of the training process, the server selects clients to participate in the round and dispatches the model to these selected clients. These clients subsequently conduct local training on the ML model and then return the trained model to the server. The server aggregates these models based on specified rules to update the global model. If the training’s end condition is met, the learning process concludes; otherwise, training persists.
2.1.2 Categories of FL Systems
FL can be categorized into different types based on the distribution of data and the devices involved, including horizontal FL(HFL), vertical FL(VFL), and federated transfer learning(FTL) [23, 24]. The categories of FL systems are presented in Fig. 3.
Figure 3: Categories of FL systems. Based on data distribution, FL can be classified into horizontal FL, vertical FL, and federated transfer learning. Additionally, based on devices involved, FL can be classified into cross-device FL and cross-silo FL.
1) Based on data distribution.
(1) Horizontal federated learning: also known as sample-based FL, refers to a scenario in which different computing nodes possess distinct sample IDs but share the same feature spaces. In this approach, data is horizontally partitioned, meaning each client holds different samples of the same features. For instance, consider two general hospitals located in different geographical regions, they may have patient databases with the same set of features but with different sets of patients.
(2) Vertical federated learning: often referred to as feature-based FL, applies to situations where feature spaces on different computing nodes are dissimilar, but sample IDs are identical. It involves data that is vertically partitioned, where each client possesses different attributes or features of the same set of examples. Imagine two specialized hospitals located in similar geographical areas, each specializing in different types of treated diseases. While they may share the same set of patients, their patient databases could have distinct feature spaces. Unlike horizontal FL, vertical FL necessitates ID alignment before the training process can commence.
(3) Federated transfer learning: Federated transfer learning addresses scenarios where data owners possess non-overlapping feature spaces and sample IDs. This category deals with the transfer of knowledge or models across different but related tasks in FL settings. For example, consider two geographically distant specialized hospitals, each treating different types of diseases. They may have databases with different sample IDs and feature spaces.
In our investigation of BC-FL systems, we observe that horizontal FL is the most commonly used approach. Only a limited number of papers employ vertical FL [25, 26] or federated transfer learning (FTL) [27]. Consequently, unless explicitly stated, when we refer to FL in this paper, we are primarily referring to horizontal FL by default.
2) Based on Devices Involved.
(1) Cross-device federated learning: Cross-device FL involves an extensive array of mobile or IoT devices utilized as training devices. In an industrial-grade cross-device setting, data is distributed among a substantial number of devices, sometimes reaching as high as $10^{10}$. This type of federated learning involves multiple devices or edge devices participating in the collaborative model training process. However, it is important to note that these devices are not dedicated computing nodes and are often constrained by their computational power and network conditions. Only a subset of the training devices can be used during each training session, and a fraction of them may experience interruptions during training (approximately 5%) [23]. Cross-device FL is typically implemented using horizontal FL, which is also the pioneering FL method. For example, Google employs this approach to train Android app functions [28, 29], while Apple utilizes it to train speech recognition systems [30].
(2) Cross-Silo Federated Learning:** Cross-silo FL involves relatively professional clients maintained by dedicated organizations, typically with fewer than 100 clients. Cross-silo FL focuses on collaboration between different data silos or entities, each with its own set of clients or devices. Training devices in cross-silo FL are typically online continuously, boasting substantial computational power and abundant communication resources. They generally do not experience interruptions during the training process. Cross-silo FL can be trained using various methods, including horizontal FL, vertical FL, or a hybrid approach, and it is frequently employed in multi-institutional federated learning scenarios, such as specific medical domains [31].
@@ 2.2 Blockchain
This section seeks to offer readers a concise overview of the essential elements of blockchain technology, encompassing its classification and consensus algorithms. Our objective is to provide readers with a broad understanding of the blockchain landscape.
2.2.1 Blockchain Basic Content
Blockchain is a distributed ledger system renowned for its decentralization [32]. In the blockchain ecosystem, no single entity can control its operations, making it resilient to the presence of potentially malicious nodes within the system [33]. By harnessing established cryptographic principles, blockchain has devised a suite of operations that necessitate consensus among multiple nodes. This innovation empowers entities with mutual distrust to collaboratively store data using blockchain technology. The fundamental structure of a blockchain is depicted in Figure 4, comprising the block header and the block body.
Figure 4: The structure of a blockchain consists of two primary components: the block header and the block body. The block header serves as a metadata repository, recording essential information about the block. It generally stores information about the block, such as the block number, timestamp, and hash of the previous block. Meanwhile, the block body encompasses a series of transaction records, containing all relevant transactional information associated with the current block. One noteworthy feature is that each block stores hash information from the previous block, ensuring the immutable and unchangeable nature of the blockchain.
2.2.2 Consensus Mechanism
Blockchain operates as a decentralized system, devoid of a central authoritative node to coordinate its functions. Instead, participants within the blockchain ecosystem rely on pre-established consensus algorithms. These consensus algorithms serve as a set of rules and algorithms that ascertain the state of the blockchain, making them indispensable components of the blockchain system. Currently, there are two primary categories of consensus mechanisms employed for generating new blocks, essentially recording data into the blockchain ledger. These mechanisms are categorized as proof-based and committee-based mechanisms, as discussed in a survey by Xu et al. [34].
(1) Proof-Based Consensus Mechanisms. Proof-based consensus mechanisms accord higher priority for bookkeeping to participants possessing more resources. Notable examples of such mechanisms include Proof of Work (PoW) and Proof of Stake (PoS) [35]. PoW assigns higher accounting priority to nodes with substantial computing resources and is widely adopted in contemporary BC-FL systems. In PoW, competing nodes solve a cryptographic puzzle, with the first node to discover the solution earning the privilege to add a block to the blockchain. In contrast, PoS prioritizes nodes based on their holdings of tokens (Stake) and offers a more energy-efficient consensus mechanism compared to PoW. PoS selects a leader through an encrypted random algorithm to generate a new block, where the likelihood of a participant being chosen depends on its token holdings. Generally, proof-based consensus mechanisms exhibit greater resistance to malicious nodes, rendering them a prevalent choice in public blockchain networks.
(2) Committee-Based Consensus Mechanisms. Committee-based consensus mechanisms typically involve an institution or organization tasked with overseeing the management of participants’ identities. The fundamental design philosophy behind these mechanisms is that consensus is achieved when the number of votes in favor of a proposal surpasses a specific threshold. Notable examples of this category of consensus mechanisms include Raft [36] and PBFT [37]. Committee-based consensus mechanisms generally deliver higher throughput compared to proof-based consensus mechanisms but are typically employed in permissioned blockchain networks. However, this protocol presents challenges in large-scale networks due to its communication-intensive nature.
@@ 3 Blockchain-empowered Federated Learning
3.1 Decentralization Architecture
In traditional FL, the parameter server acts as the central point of communication for clients to download the current global model, train on it, and upload the updated local model. This architecture is inherently centralized, with the parameter server acting as the hub of all communication between clients. The client needs to maintain continuous communication with the sole parameter server throughout the entire FL training process. However, such an architecture exhibits a single point of failure risk - any failure of the parameter server, whether due to an attack or technical issue, would suspend the entire learning process and potentially lead to data loss. Additionally, there is always a risk that the parameter server itself may not be completely trustworthy, which can lead to incorrect aggregation of models and undermine the effectiveness of the FL system.

### v2

@@ 1 Introduction
Artificial Intelligence (AI) technologies drive the Fourth Industrial Revolution, with user data being essential for training diverse Machine Learning (ML) models [1]. Training high-quality ML models often involves a centralized approach, necessitating internal storage of user data. This raises privacy concerns [2, 3, 4] and highlights the need for stringent privacy protections [5]. In recent years, regions such as the European Union [6, 7], the United States [8], and Singapore [5] have enacted relevant laws and regulations to govern the use of personal data, enhancing privacy protection but potentially hindering the utilization of high-quality data.
Federated Learning (FL) is a privacy-preserving distributed machine learning paradigm that balances user data protection and effective utilization [9, 10, 11]. FL involves training local models on user devices and aggregating these local models into a global model on a server without requiring users to upload their data, addressing the aforementioned privacy concerns. Initially applied to training Gboard [12], FL has proven successful. Its potential extends beyond this, as it can also address the issue of data silos. Data silos refer to the isolated or dispersed nature of data, making access to this data extremely challenging [13]. One cause of data silos is the reluctance of organizations to share data due to privacy or competitive concerns. For instance, due to privacy protection, hospitals may be unwilling to share patient data [14]. In summary, the judicious use of FL can break down data barriers, leading to its widespread application in healthcare [15, 16], finance [17, 18], industry [19, 20] and so on.
While privacy protection and data utilization benefits have popularized FL across industry and academia, they also introduce specific challenges. First, there is a lack of trust among nodes within the FL system [21, 22]. Nodes may worry that their training contributions will be intentionally tampered with or miscalculated, damaging their reputation and deserved rewards. Second, FL systems are vulnerable to attacks from malicious nodes [23, 24]. Malicious users may intentionally provide incorrect information to prevent model convergence and disrupt model training, while malicious servers can recover users’ training data from the uploaded models. Third, FL is prone to single point of failure issues [25]. In traditional FL architectures, the central server is responsible for aggregating and updating global model parameters. If the central server is attacked or fails, the entire system’s operation is severely affected, leading to interruptions in the training process, data loss, and irrecoverable model states.
Blockchain is essentially a distributed ledger, and its successful application in cryptocurrencies demonstrates its potential to build trust, security, and transparency [26, 27, 28]. Consequently, numerous studies have integrated blockchain with FL systems to enhance functionality, creating blockchain-empowered FL (BC-FL) systems. Analyzing existing BC-FL literature, we find that blockchain’s enhancement of different aspects of FL originates from its distinct properties. First, blockchain’s transparency and immutability can alleviate the lack of trust among nodes within the FL system. By recording data requiring consensus in the FL system on the blockchain, these data cannot be tampered with by malicious nodes, enhancing trust relationships. Second, through cross-validation of blockchain nodes and other mechanisms, the resistance of the FL system to malicious nodes is improved. Finally, blockchain can replace the centralized server to avoid single point of failure issues. By designing a reasonable consensus mechanism, suitable clients can be selected to undertake model aggregation tasks in each communication round. With the advent of blockchain 2.0, users can develop smart contracts running on the blockchain, endowing BC-FL with greater scalability for automatically running various algorithms [29].
Figure 1: Main scope of this survey. We begin by exploring the characteristics of blockchain and investigate its enhancement of Federated Learning systems. Next, we discuss the additional challenges introduced by using blockchain in FL systems and review existing solutions. Finally, we outline future research directions for Blockchain-empowered Federated Learning systems.
The introduction of blockchain has further driven the development of FL, but blockchain is not a panacea for FL. Our research indicates that blockchain integration poses challenges related to runtime efficiency and storage capacity. First, the consensus mechanism of blockchain adds communication and computation overhead to the BC-FL system. Second, due to the distributed storage nature of blockchain, full nodes need to back up the entire blockchain data. Additionally, the introduction of blockchain can also bring additional security issues, such as Sybil attacks [30].
Currently, several surveys on BC-FL systems have been published. Some focus on the integration of BC-FL with other fields, such as the Internet of Things, drones, and healthcare. These studies emphasize the specific applications of BC-FL systems rather than their commonalities. Other surveys investigate BC-FL systems in general. Qu et al. conducted a detailed study on the performance of decentralization, attack resistance, and incentive mechanisms in BC-FL systems, and surveyed the system architecture forms of BC-FL. However, they did not investigate transparent reputation mechanisms in BC-FL and thoroughly analyze why blockchain can enhance FL systems, merely classifying the functions of BC-FL. Zhu et al. divided BC-FL system models into three categories and surveyed real-world applications of BC-FL. However, they lacked a comprehensive investigation of single point of failure, reputation mechanisms, security, and privacy issues. Additionally, the aforementioned surveys neglected to conduct a detailed investigation into the negative effects blockchain can bring to BC-FL systems. Sameera et al. summarized the general architecture of BC-FL and detailed how BC-FL addresses security and privacy threats. However, they lacked in-depth research on BC-FL’s reputation mechanisms, incentive mechanisms, and system efficiency and storage issues. We believe that the various enhancements and potential challenges blockchain brings to FL systems stem from certain properties or functionalities of blockchain. Meanwhile, some properties of blockchain can play different roles in FL systems depending on their application. The contributions of this work are as follows:
  • Starting from the characteristics and functionalities of blockchain, we introduce how blockchain enhances FL systems in terms of decentralization, reputation mechanisms, incentive mechanisms, and security.
  • We comprehensively investigate the additional challenges that arise from using blockchain in FL systems, the reasons behind these challenges, and existing solutions.
  • We summarize future research directions for blockchain-based FL systems based on existing research.
Investigating how blockchain enhances FL systems from the perspective of blockchain’s characteristics is an unaddressed area in existing surveys. This survey can complement similar recent surveys, filling a gap in the research on BC-FL. This survey.’s main scope is illustrated in Fig. 1. Section 2 provides relevant background knowledge on FL, blockchain. Section 3 details general BC-FL architecture and how blockchain enhances FL systems. Section 4 elaborates on the challenges of using blockchain in FL systems and existing solutions. Section 5 we point out future research directions for BC-FL systems. Section 6 concludes the paper.
Figure 2: Categories of FL systems.
@@ 2.1 Federated Learning
FL is a privacy-preserving distributed machine learning paradigm proposed by Google [9]. This paradigm involves a network of multiple participants (clients) alongside a central server. The clients are tasked with developing local models, which are then consolidated by the server into a unified, global model [31, 32, 33]. This structure allows participants considerable autonomy, enabling them to contribute to the FL framework without disclosing their data to any FL node. Participation in FL training remains at the discretion of the users. Fig. 3 visually represents the standard FL training methodology. The task initiator selects an FL server to publish the training task. Subsequently, clients related to the training task join the FL training, and the server initializes the global model. In each communication round, the server selects clients to participate in that round of training and distributes the global model to these clients. The clients then use their local datasets to train the global model, resulting in local models, which they send back to the server. The server aggregates these local models into a new global model according to certain rules. The FL process stops when the training termination condition is met; otherwise, training continues.
FL generally has two classification methods, as depicted in Fig. 2. Based on the sample ID and feature distribution of local datasets, FL systems can be divided into Horizontal FL (HFL), Vertical FL (VFL), and Federated Transfer Learning (FTL). In HFL, users have different sample IDs but identical data features. In VFL, users have identical sample IDs but different data features. In FTL, both the sample IDs and data features differ. Generally, current research on BC-FL is predominantly based on HFL, with only a few studies on VFL [34, 35] and FTL [36]. Despite the datasets’ different characteristics, blockchain’s role does not fundamentally differ.
Figure 3: Standard federated learning (FL) training methodology.
Moreover, based on the different types of user devices involved in training, FL can be divided into cross-device FL and cross-silo FL. Cross-device FL involves an extensive array of mobile and IoT devices, potentially numbering up to $10^{10}$, often with poor device performance and network conditions [37, 38]. In cross-silo FL systems, the participating clients are typically professional computing nodes maintained by specialized institutions, and the number of clients is generally fewer than 100 [39, 40].
@@ 2.2 Blockchain
Blockchain is a distributed ledger system designed to securely, transparently, and immutably record data in a decentralized manner [41, 42]. It is resilient to the presence of some malicious nodes and is resistant to control by any single entity [43, 44]. A blockchain is essentially composed of interconnected blocks arranged sequentially in a chronological chain. This structure is illustrated in the Fig. 4. Each block includes a header and a body. The block header typically contains metadata such as the block number, timestamp, and the hash of the previous blocks. The specific parameters stored in the header can vary depending on the blockchain system. The block body stores specific information, like transactions.
Blockchain can be classified into two main types based on node participation restrictions: permissionless blockchain and permissioned blockchain. The permissionless blockchain allows unrestricted node participation in the system’s operations, exemplified by Bitcoin and Ethereum. These blockchains operate without the need for approvals or authorizations from central authorities. In contrast, the permissioned blockchain is managed by a specific organization, and only authorized nodes can access the system. It is suitable for data privacy and security applications, such as financial and government information management.
The autonomous operation of blockchains in the absence of central oversight relies on consensus mechanisms. These algorithms define the state of the blockchain and are vital for its functionality. Broadly, consensus mechanisms are categorized into proof-based and committee-based systems [26]. Proof-based systems prioritize nodes based on specific resource possession; for instance, Proof of Work (PoW) and Proof of Stake (PoS) are notable examples [45]. In PoW, nodes compete to solve complex puzzles, with computational prowess conferring a higher accounting priority. PoS, on the other hand, employs an encrypted random selection algorithm to appoint a leader for block creation, with a node’s selection likelihood tied to its token holdings. Proof-based consensus mechanisms are predominantly utilized in permissionless blockchains due to their robust defence against malicious nodes. Conversely, committee-based mechanisms utilize a voting process where consensus on a new block’s addition is reached through a predetermined number of affirmative votes. Protocols such as Raft [46] and PBFT [47] exemplify committee-based mechanisms. Committee-based mechanisms generally offer higher throughput than proof-based mechanisms and are often used in permissioned blockchains. However, they can be communication-intensive, posing challenges in large-scale networks.
Figure 4: The structure of a blockchain.
@@ 3 Blockchain-empowered Federated Learning
This section introduces how blockchain technology enhances the FL system. These enhancements can be categorized into four aspects: decentralization, reputation evaluation mechanism, incentive evaluation mechanism, and security. Each of these four aspects will be discussed separately.
3.1 Decentralization
FL traditionally relies on a central parameter server, where clients must continuously communicate with a single FL server. This centralized structure poses significant risks, such as single points of failure and potential malicious server behaviour. Furthermore, node reputation information is solely managed by the server, which is not ideal for developing an open FL ecosystem. Blockchain’s decentralization is a core feature that fundamentally addresses these issues and provides a new architectural approach to FL systems. Decentralization can enhance the security, transparency, and reliability of FL systems, with these benefits manifesting in various facets of blockchain’s advantages for FL.
Our review of existing literature identifies several factors influencing the decentralization of BC-FL, including system architecture, consensus mechanisms, and smart contracts. The system architecture determines which nodes maintain the blockchain, while the consensus mechanism dictates which nodes have the right to manage the system. Smart contracts can automate various algorithms within the BC-FL system, offering greater scalability. Table 1 compiles representative BC-FL systems, detailing their use of smart contracts, architecture, consensus mechanisms, and experimental platforms.
Table 1: BC-FL Systems Based on Blockchain and Federated Learning.
Ref. Smart contract Architecture Consensus algorithm Platform
Abdel [48] $\checkmark$ Complete Algorand Other
Fang [49] $\times$ Partial Algorand Other
Feng [50] $\checkmark$ Partial PBFT/Raft Hyperledger Fabric
Guo [51] $\checkmark$ Partial PBFT Hyperledger Fabric
Jiang [52] $\times$ Complete DPoS Other
Liu [53] $\times$ Partial PoW+PoA Other
Lu [54] $\times$ Complete DPoS Other
Nguyen [55] $\times$ Partial PoR Other
Nguyen [56] $\times$ Partial PoW Other
Qi [57] $\checkmark$ Partial - Hyperledger Fabric
Qi [58] $\checkmark$ Complete Modified PBFT Ethereum
Qu [59] $\times$ Complete PoW Other
Rehman [60] $\checkmark$ Complete - Ethereum
Wu [61] $\times$ Complete PoW Other
Xu [62] $\times$ Complete - Other
Xu [63] $\checkmark$ Complete - Other
Zhang [64] $\times$ Partial PoW Other
Zhao [65] $\times$ Partial PBFT Other
Wang [66] $\times$ Complete - Other
Huang [66] $\checkmark$ Partial Raft, HotStuff FISCO
Ouyang [67] $\checkmark$ Partial PoS Ethereum
Yuan [68] $\checkmark$ Partial Raft, Modified DAG Hyperledger, DAG
aloqaily [69] $\times$ Partial - Other
Mu [70] $\times$ Complete - Other
Wahrstatter [71] $\times$ Complete PoS Ethereum
Architecture. BC-FL systems can be categorized by their degree of decentralization: complete and partial. In the completely decentralized BC-FL, all nodes are eligible to participate in the consensus process of the blockchain. Fig. 5(a) shows its general system architecture. This approach demands high computational and storage capacities from all nodes. Conversely, partially decentralized BC-FL involves only a subset of nodes running the blockchain, while others focus solely on FL training. Fig. 5(b) shows its general system architecture. The selected nodes that operate the blockchain system are known as super nodes and typically have stronger computing power and better communication conditions. This approach sacrifices some transparency for increased efficiency.
Consensus Mechanism. A considerable portion of the work adopts common blockchain consensus mechanisms such as PoW and PBFT. PoW involves blockchain nodes competing to solve a mathematical problem, with the first solver aggregating models and training information into a new block. Other nodes then verify the block’s correctness, and upon majority approval, it is added to the blockchain. In PBFT, a set of consensus nodes is chosen within the BC-FL system, from which a leader node aggregates the model and generates a new block. Other nodes in the set verify the leader’s block. Some work has specifically developed consensus mechanisms for BC-FL, such as Proof of Reputation [55]. These custom consensus mechanisms are usually designed to enhance FL functionality or mitigate the disadvantages of blockchain, which will be elaborated on in the subsequent sections.
Smart Contracts. Smart contracts significantly enhance the scalability of BC-FL systems. For instance, model aggregation can be executed via smart contracts, increasing transparency. Additionally, smart contracts can deploy algorithms for detecting and handling malicious nodes, thereby improving system efficiency. They can also manage node reputation evaluations and incentive algorithms, further enhancing system transparency.